A body starts from rest and accelerates uniformly at a rate of 5m/s2. Calculate its velocity after 90m
1
Expert's answer
2019-03-11T11:55:41-0400

The distance traveled can be expressed in terms of initial velocity, final velocity and acceleration this way:


"d=\frac{v_\text{f}^2-v_\text{i}^2}{2a},"

but since the body accelerates from rest, its initial velocity is zero and the final velocity equals


"v_\text{f}=\sqrt{2da}=\sqrt{2\cdot 90\cdot 5}=30\text{ m/s}."